OpenStudy (anonymous):

Use an appropriate test to determine whether the series is absolutely convergent, conditionally convergent or divergent

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty} (-1)^{i+1}\frac{ \sqrt(i) }{ i+3 }\]

OpenStudy (anonymous):

I did the LCT and my test series is \[\sum_{i=1}^{\infty} \frac{ i^{1/2} }{ i }\]

OpenStudy (austinl):

Would ratio test be appropriate? Sorry, I'm a little rusty.

OpenStudy (anonymous):

Oh maybe let me try

OpenStudy (austinl):

Actually, try the alternating series test.

OpenStudy (anonymous):

Okay i did that and I got that it converges conditionally. Do you think that's right

OpenStudy (anonymous):

Yay!!!! Thanks!